OpenStudy (anonymous):

Which of the following is a solution of x2 - 8x = -27

OpenStudy (jojo4eva):

ight what are the solutions

OpenStudy (anonymous):

\[A)-4 - \sqrt{11} B) 4+i \sqrt{11} C) -4-i \sqrt{44} D) 4+i \sqrt{44}\]